# Enhancing Interpretability in Deep Reinforcement Learning through Semantic Clustering

### v4

@@ Abstract
In this paper, we explore semantic clustering properties of deep reinforcement learning (DRL) to improve its interpretability and deepen our understanding of its internal semantic organization. In this context, semantic clustering refers to the ability of neural networks to cluster inputs based on their semantic similarity in the internal space. We propose a DRL architecture that incorporates a novel semantic clustering module that combines feature dimensionality reduction with online clustering. This module integrates seamlessly into the DRL training pipeline, addressing the instability of t-SNE and eliminating the need for extensive manual annotation inherent to prior semantic analysis methods. We experimentally validate the effectiveness of the proposed module and demonstrate its ability to reveal semantic clustering properties within DRL. Furthermore, we introduce new analytical methods based on these properties to provide insights into the hierarchical structure of policies and semantic organization within the feature space.
@@ 1 Introduction
Deep reinforcement learning (DRL) has been widely applied in domains such as robotics, autonomous systems, game playing, and healthcare, due to its ability to solve complex decision-making tasks [6, 13, 24]. However, the black-box nature of DRL models obscures the decision-making process, potentially leading to unforeseen risks. In this study, we explore semantic clustering properties to improve the interpretability of DRL models. We use the term semantic clustering to refer to the process of grouping information based on underlying semantic similarities in the feature space. Studying semantic clustering in DRL helps reveal the model’s internal knowledge structure and semantic relationships between states, enhancing the interpretability and transparency of DRL models.
Although semantic clustering has been thoroughly investigated in natural language processing (NLP) [21, 20] and computer vision (CV) [14, 22, 19, 34], it remains underexplored in DRL, due to the complexity introduced by temporal dynamics and the absence of direct supervised signals. The sequential nature of decisions in DRL further complicates the task of capturing evolving semantics. Early work introduced external constraints—e.g., bisimulation [9, 33] and contrastive learning [8, 1, 11]—to shape feature spaces conducive to semantic clustering. In contrast, we focus on investigating whether DRL can intrinsically exhibit semantic clustering capabilities.
Mnih et al. [17] and Zahavy et al. [32] analyzed the semantic distribution of the DRL feature space for Atari games using t-distributed Stochastic Neighbor Embedding (t-SNE) [27]. However, these studies are limited in multiple ways: (i) they are limited to a small set of Atari games with fixed scenes, making it difficult to distinguish whether clustering arises from pixel similarity or semantic understanding, (ii) Zahavy et al. [32] manually define features for specific games, imposing substantial human effort, and (iii) both studies rely on t-SNE visualization for semantic analysis, which tends to produce unstable results and lacks an automated clustering mechanism. Thus, these approaches require significant manual effort for feature space annotation and analysis, hindering comprehensive semantic analysis and integration into downstream tasks.
@@
  • We comprehensively explore the semantic clustering properties of DRL, advancing the understanding of the black-box decision-making processes. Unlike prior work that uses fixed-scene Atari games, we use Procgen  [^1] Detailed environment instructions are available here, in Appendix A of Cobbe et al. [4], and in the repository. [4], which offers rich semantic diversity and dynamic, procedurally generated environments.
  • We introduce a novel architecture that addresses limitations in previous t-SNE-based semantic analysis methods by providing a more stable and effective approach for studying semantic properties in DRL.
@@ 2 Related Work
Semantic Clustering in DRL Mnih et al. [17] and Zahavy et al. [32] have previously explored visualizing the DRL feature space using t-SNE. In these studies, t-SNE visualizations show that features of states with close pixel distances tend to cluster together. However, due to the fixed nature of the scenarios they used (Atari games), semantic clustering could not be conclusively verified. This limitation motivates our use of Procgen to validate our approach.
Interpretability of DRL DRL interpretability research often focuses on video games due to their controlled environments and clear rules, which make analyzing decision-making processes easier. PW-Net [10] enhances interpretability by using human-friendly prototypes to explain the model’s decision-making. DIGR [30] improves interpretability by generating saliency maps that highlight the most relevant features influencing the agent’s decisions. Concept policy models integrate expert knowledge into multi-agent RL, enabling real-time intervention and interpretation of agent behavior [31]. MENS-DT-RL [5] applies decision trees to provide a rule-based explanation of the learning process. Furthermore, attention mechanisms and symbolic reasoning frameworks have also been applied to enhance interpretability [25, 18, 15]. Our work explores how DRL models internally organize information, offering new insights into the structure of learned representations.
@@ 3 Method
The VQ-VAE [26] workflow begins with an encoder network $\hat{E}$ that maps an input $\mathbf{x}$ to a latent representation $\hat{E}(\mathbf{x})$. This representation is then quantized by mapping it to the nearest embedding in a codebook $\left\{\mathbf{e}_{k}|k\in\left\{1,2,\dots,K\right\}\right\}$. The quantized representation is then passed into a decoder network $\hat{D}$ to reconstruct the input $\mathbf{x}$. The loss function for VQ-VAE is defined as:
@@
In this paper, we modify VQ-VAE to achieve the following objectives: (i) assign features to the nearest VQ embedding for clustering, (ii) seamlessly integrate with DRL training, enabling simultaneous clustering and policy learning, and (iii) enhance clustering and interpretability through joint training with additional losses. Further details are provided in subsection 3.2.
Figure 1: Overview of our architecture. The upper segment represents the classic DRL training pipeline, while the lower segment introduces the semantic clustering module. The Feature Dimensionality Reduction (FDR) net reduces the dimensionality of state features, resulting in FDR features, which the vector quantizer then processes to generate discrete VQ codes (denoted $k$), which represent states associated with clusters, along with the closest VQ embeddings. Subsequently, $k$ is integrated into the state feature by element-wise addition after being expanded to match the state feature dimensions, enabling conditional policy training that better supports the integration of downstream tasks.
@@ 3.1 Semantic Clustering Module
To overcome the limitations of previous t-SNE-based semantic analyses (see section 1), we propose a novel semantic clustering module, which includes dimensionality reduction and online clustering.
Dimensionality Reduction Given the complexity of states in most DRL applications, their features are often high-dimensional. For example, Mnih et al. [17] use a 512-dimensional feature vector when training DQN on Atari games. Clustering high-dimensional features is challenging due to "the curse of dimensionality" [3, 2]. To mitigate these issues, we reduce feature dimensionality before clustering, resulting in more robust clustering outcomes. This not only simplifies the clustering process but also enables human-interpretable visualizations, typically in 2D.
The instability of t-SNE arises from its non-convex objective function, making it highly sensitive to initialization and leading to varied and unstable visualization outcomes [27, 28]. To overcome these challenges, we propose the Feature Dimensionality Reduction (FDR) network. This network remaps high-dimensional features to 2D using policy training data for online training, ensuring stable and efficient mappings after training. The FDR network’s loss function is designed to preserve the consistency of distance relationships between high-dimensional and 2D feature spaces, measured by pairwise similarities as described in subsection 3.2.
Online Clustering To reduce the need for extensive manual annotation from previous work and to facilitate the integration of downstream tasks, we propose using online clustering to segment the feature space and assist in semantic analysis. We modify the VQ-VAE framework to achieve online clustering, as detailed in subsection 3.2.
@@ 3.2 Loss Function Design
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{DRL}}+f_{\text{control}}\left(w_{\text{FDR}}\mathcal{L}_{\text{FDR}}+w_{\text{VQ-VAE}}\mathcal{L}^{\prime}_{\text{VQ-VAE}}\right).$ (2)
The DRL loss function $\mathcal{L}_{\text{DRL}}$ comes from the original DRL model. $w_{\text{FDR}}$ and $w_{\text{VQ-VAE}}$ are the weights of the FDR loss ($\mathcal{L}_{\text{FDR}}$) and the modified VQ-VAE loss ($\mathcal{L}^{\prime}_{\text{VQ-VAE}}$), respectively. $f_{\text{control}}$ represents the control factor. We explain each of these components below.
@@
The pairwise similarities of state features $p_{ij}$ is given by
@@
where $f$ is the feature extractor, $\mathbf{s}_{i}$ is the $i^{\text{th}}$ state in a batch, and $\alpha$ is the degree of freedom. The pairwise similarities for FDR features, $q_{ij}$ are computed using the same expression as (3), but with $f$ replaced by $g\circ f$, where $g$ is the FDR net. In contrast to other deep clustering studies, e.g., Xie et al. [29] and Li et al. [12], the same degree of freedom $\alpha$ is selected for both high- and low-dimensional similarities, ensuring that the original distance relationship between features is maintained in the low-dimensional space.
@@
Control Factor Since effective semantic clustering relies on a clear and distinguishable semantic distribution that is often difficult to achieve in the early stages of training, we propose an adaptive control factor ($f_{\text{control}}$) strategy updated according to training performance (see Appendix A).
Improved Clustering Our loss design not only achieves dimensionality reduction and clustering but also enhances clustering properties, making the states within each cluster more compact and the cluster boundaries more separable. This improvement is crucial for clearly distinguishing the semantics of states at the cluster boundaries, further enhancing the model’s interpretability. We demonstrate the improved clustering in section 4 and provide more evidence of this enhanced clustering property and the intrinsic nature of semantic clustering in DRL in subsection C.2.
This enhanced clustering arises from the fact that although the stop-gradient operation ($sg$) in $\mathcal{L}^{\prime}_{\text{VQ-VAE}}$ prevents FDR features from moving towards the closest VQ embeddings, when the FDR features become denser under the influence of $\mathcal{L}_{\text{FDR}}$, it indirectly reduces $\mathcal{L}^{\prime}_{\text{VQ-VAE}}$ and, consequently, $\mathcal{L}_{\text{total}}$. In other words, training with both $\mathcal{L}_{\text{FDR}}$ and $\mathcal{L}^{\prime}_{\text{VQ-VAE}}$ decreases the sum of distances between FDR features and their corresponding VQ embeddings within the same cluster. Additionally, since $\mathcal{L}_{\text{FDR}}$ aims to bring $p$ and $q$ closer, the improved cluster density in the FDR feature space is also transferred to the state feature space.
@@
2 for each training iteration $i=1,2,\dots$ do
@@
6
       $\mathcal{L}_{\text{PPO}}\leftarrow\mathcal{L}_{\text{policy}}+w_{\text{value}}\cdot\mathcal{L}_{\text{value}}+w_{\text{entropy}}\cdot\mathcal{L}_{\text{entropy}}$;
@@
       $\mathcal{L}_{\text{SCM}}\leftarrow w_{\text{FDR}}\cdot\mathcal{L}_{\text{FDR}}+w_{\text{VQ-VAE}}\cdot\mathcal{L}^{\prime}_{\text{VQ-VAE}}$;
@@
       $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{PPO}}+f_{\text{control}}\cdot\mathcal{L}_{\text{SCM}}$;
@@
Advantages of Online Training Online training offers several key advantages: (i) It enhances clustering by incorporating the training of $\mathcal{L}_{\text{total}}$. (ii) Training the VQ code $k$ with a latent-conditioned policy $\pi\left(a|\mathbf{s},k\right)$ (where $a$ is the action) supports the extension to downstream tasks, such as macro action selection in hierarchical learning. (iii) Online training improves memory efficiency by eliminating the need to store a large number of states during model training.
@@ 4 Simulations
In this work, we primarily study the intrinsic characteristics and black-box decision-making of DRL, and address the instability of t-SNE visualizations used in prior studies. Therefore, this section aims to: (i) compare t-SNE to validate the stability and effectiveness of the proposed clustering method; (ii) assess the semantic clustering properties of DRL to improve interpretability; and (iii) introduce new methods to analyze policies and internal model characteristics, identifying issues in DRL decision-making. For performance analysis, please refer to subsection C.1 and Appendix D, which demonstrate that the integration of our module has minimal impact on performance.
@@ 4.1 Clustering Effectiveness Evaluation
We demonstrate the clustering effectiveness of our proposed approach using the CoinRun game from Procgen as an example. Similar results can be easily extended to other games using the code and checkpoints provided in the supplementary material. We use a trained model to collect states, where the agent selects actions randomly with a probability of 0.2 and follows the trained policy with a probability of 0.8 to ensure diverse state coverage. States are sampled with a probability of 0.8, and 64 parallel environments collect states over 500 steps, resulting in $\sim$25,000 states for visualization.
Cluster Separation and Improved Clustering The t-SNE visualization of PPO (2(a)), spreads features across the space without forming clear clusters, limiting its utility for clustering analysis and requiring detailed manual examination of certain areas, as in previous studies. In contrast, the t-SNE visualization of our method (2(b)), reveals numerous distinct, small clusters. States within each of these clusters originate from the same semantic group identified by our method. This dispersion into multiple smaller clusters is due to t-SNE’s focus on local structures and its tendency to avoid crowding, causing complete semantic clusters to scatter. The visualization in the FDR space (2(d)), displays clear and separate complete clusters, which are identified by VQ codes.
@@
Sensitivity to Random Seed While the t-SNE representation is sensitive to randomness, as demonstrated by the significant difference between Figures 2(c) and 2(f), the FDR space’s mapping remains unchanged even when the random seed is altered, maintaining the distribution in 2(e). t-SNE’s randomness primarily stems from its random initialization and non-convex optimization process, leading to different visualizations with different random seeds. In contrast, our model produces a stable feature mapping after training, which does not vary with random seeds.

### v5

@@ Abstract
In this paper, we explore semantic clustering properties of deep reinforcement learning (DRL) to improve its interpretability and deepen our understanding of its internal semantic organization. In this context, semantic clustering refers to the ability of neural networks to cluster inputs based on their semantic similarity in the feature space. We propose a DRL architecture that incorporates a novel semantic clustering module that combines feature dimensionality reduction with online clustering. This module integrates seamlessly into the DRL training pipeline, addressing the instability of t-SNE and eliminating the need for extensive manual annotation inherent to prior semantic analysis methods. We experimentally validate the effectiveness of the proposed module and demonstrate its ability to reveal semantic clustering properties within DRL. Furthermore, we introduce new analytical methods based on these properties to provide insights into the hierarchical structure of policies and semantic organization within the feature space. Our code is available at https://github.com/ualiangzhang/semantic_rl.
@@ 1 Introduction
Deep reinforcement learning (DRL) has been widely applied in domains such as robotics, autonomous systems, game playing, and healthcare, due to its ability to solve complex decision-making tasks [1, 2, 3]. However, the black-box nature of DRL models obscures the decision-making process, potentially leading to unforeseen risks. In this study, we explore semantic clustering properties to improve the interpretability of DRL models. We use the term semantic clustering to refer to the process of grouping states eliciting similar agent behaviors under comparable environmental contexts (e.g., approaching a target, jumping to a higher platform in Procgen). Studying semantic clustering in DRL helps reveal the model’s internal knowledge structure and semantic relationships between states, enhancing the interpretability and transparency of DRL models.
Although semantic clustering has been thoroughly investigated in natural language processing (NLP) [4, 5] and computer vision (CV) [6, 7, 8, 9], it remains underexplored in DRL due to the complexity introduced by temporal dynamics and the absence of direct supervised signals. The sequential nature of decisions in DRL further complicates the task of capturing evolving semantics. Early work introduced external constraints—e.g., bisimulation [10, 11] and contrastive learning [12, 13, 14, 15]—to shape feature spaces conducive to semantic clustering. In contrast, we focus on investigating whether DRL can intrinsically exhibit semantic clustering capabilities.
[16] and [17] analyzed the semantic distribution of the DRL feature space for Atari games using t-distributed Stochastic Neighbor Embedding (t-SNE) [18]. However, these studies are limited in multiple ways: (i) they are limited to a small set of Atari games with fixed scenes, making it difficult to distinguish whether clustering arises from pixel similarity or semantic understanding, (ii) [17] manually define features for specific games, imposing substantial human effort, and (iii) both studies rely on t-SNE visualization for semantic analysis, which tends to produce unstable results and lacks an automated clustering mechanism. Thus, these approaches require significant manual effort for feature space annotation and analysis, hindering comprehensive semantic analysis and integration into downstream tasks.
@@
  • We comprehensively explore the semantic clustering properties of DRL, advancing the understanding of the black-box decision-making processes. Unlike prior work that uses fixed-scene Atari games, we use Procgen [^1] Detailed environment instructions are available here, in Appendix A of [19], and in the repository. [19], which offers rich semantic diversity and dynamic, procedurally generated environments.
  • We introduce a novel end-to-end architecture that integrates feature dimensionality reduction with online clustering, overcoming the limitations of prior t-SNE-based analyses and providing a more stable, effective means to study semantic properties in DRL.
@@ 2 Related Work
Semantic Clustering in DRL [16] and [17] have previously explored visualizing the DRL feature space using t-SNE. In these studies, t-SNE visualizations show that features of states with close pixel distances tend to cluster together. However, due to the fixed nature of the scenarios they used (Atari games), semantic clustering could not be conclusively verified. This limitation motivates our use of Procgen to validate our approach.
Interpretability of DRL DRL interpretability research often focuses on video games due to their controlled environments and clear rules, which make analyzing decision-making processes easier. PW-Net [20] uses human-friendly prototypes to explain the model’s decision-making. DIGR [21] generates saliency maps tlat highlight the most relevant features influencing the agent’s decisions. Concept policy models integrate expert knowledge into multi-agent RL, enabling real-time intervention and interpretation of agent behavior [22]. MENS-DT-RL [23] applies decision trees to provide a rule-based explanation of the learning process. Furthermore, attention mechanisms and symbolic reasoning frameworks have also been applied to enhance interpretability [24, 25, 26]. Our work explores how DRL models internally organize information, offering new insights into the structure of learned representations.
VQ-VAE
VQ-VAE [27] is a family of generative models that combine classic VAE with discrete latent representations through a posterior parameterization. Recently, VQ-VAE has been applied to various tasks, including high-resolution image generation [28], video generation [29], and speech coding [30]. It has also been used in model-based DRL to train transition models [31, 32].
@@ 3 Method
The VQ-VAE workflow begins with an encoder network $\hat{E}$ that maps an input $\mathbf{x}$ to a latent representation $\hat{E}(\mathbf{x})$. This representation is then quantized by mapping it to the nearest embedding in a codebook $\left\{\mathbf{e}_{k}|k\in\left\{1,2,\dots,K\right\}\right\}$. The quantized representation is then passed into a decoder network $\hat{D}$ to reconstruct the input $\mathbf{x}$. The loss function for VQ-VAE is defined as:
@@
In this paper, we modify VQ-VAE to (i) assign features to the nearest VQ embedding for clustering, (ii) seamlessly integrate with DRL training, enabling simultaneous clustering and policy learning, and (iii) enhance clustering and interpretability through joint training with additional losses. Further details are provided in § 3.2.
Figure 1: Overview of our architecture. The upper segment represents the classic DRL training pipeline, while the lower segment introduces the semantic clustering module. The Feature Dimensionality Reduction (FDR) net reduces the dimensionality of state features, resulting in FDR features, which the vector quantizer then processes to generate discrete VQ codes (denoted kk)—which represent states associated with clusters—along with the closest VQ embeddings. Subsequently, $k$ is integrated into the state feature by element-wise addition after being expanded to match the state feature dimensions, enabling conditional policy training that better supports the integration of downstream tasks.
@@ 3.1 Semantic Clustering Module
To overcome the limitations of previous t-SNE-based semantic analyses (see § 1), we propose a novel semantic clustering module, which includes dimensionality reduction and online clustering.
Dimensionality Reduction Given the complexity of states in most DRL applications, their features are often high-dimensional. For example, [16] use a 512-dimensional feature vector when training DQN on Atari games. Clustering high-dimensional features is challenging due to the curse of dimensionality [33, 34]. To mitigate these issues, we reduce feature dimensionality before clustering, resulting in more robust clustering outcomes. This not only simplifies the clustering process but also enables human-interpretable visualizations, typically in 2D.
The instability of t-SNE arises from its non-convex objective function, making it highly sensitive to initialization and leading to varied and unstable visualization outcomes [18, 35]. To overcome these challenges, we propose the Feature Dimensionality Reduction (FDR) network. This network remaps high-dimensional features to 2D using policy training data for online training, ensuring stable and efficient mappings after training. The FDR network’s loss function is designed to preserve the consistency of distance relationships between high-dimensional and 2D feature spaces, measured by pairwise similarities as described in § 3.2.
Online Clustering
t-SNE–based analyses (e.g., [16, 17]) require per-state inspection and manual grouping because the plots lack clear cluster boundaries and often split semantically similar states across disconnected regions, making human curation time-consuming. To reduce such extensive annotation and facilitate downstream integration, we introduce an online clustering approach—implemented via a modified VQ-VAE—that automatically segments the feature space and supports semantic analysis. This removes manual grouping: annotators instead watch a few short clips per discovered cluster and provide a semantic summary (see Table 1), typically within $\approx$15 minutes per environment. Details of the modified VQ-VAE design are provided in § 3.2.
@@ 3.2 Loss Function Design
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{DRL}}+\lambda_{\text{ctrl}}\left(w_{\text{FDR}}\mathcal{L}_{\text{FDR}}+w_{\text{VQ-VAE}}\mathcal{L}^{\prime}_{\text{VQ-VAE}}\right).$ (2)
The DRL loss function $\mathcal{L}_{\text{DRL}}$ comes from the original DRL model. $w_{\text{FDR}}$ and $w_{\text{VQ-VAE}}$ are the weights of the FDR loss ($\mathcal{L}_{\text{FDR}}$) and the modified VQ-VAE loss ($\mathcal{L}^{\prime}_{\text{VQ-VAE}}$), respectively. $\lambda_{\text{ctrl}}$ represents the control factor. We explain each of these components below.
@@
The pairwise similarities of state features $p_{ij}$ are given by
@@
Here, $f$ is the feature extractor, $\mathbf{s}_{i}$ is the $i^{\text{th}}$ state in a batch, $\alpha$ is the Student’s-t degrees-of-freedom parameter, and $\|\cdot\|$ is the $\ell_{2}$ norm. The pairwise similarities for FDR features, $q_{ij}$ are computed using the same expression as (3), but with $f$ replaced by $g\circ f$, where $g$ is the FDR net. In contrast to other deep clustering studies, e.g., [36] and [37], the same degree of freedom $\alpha$ is selected for both high- and low-dimensional similarities, ensuring that the original distance relationship between features is maintained in the low-dimensional space.
@@
Minimizing $\mathcal{L}_{\text{FDR}}$ encourages the low‑dimensional mapping to preserve the pairwise neighborhood structure of the high‑dimensional features.
@@
Control Factor Since effective semantic clustering relies on a clear and distinguishable semantic distribution that is often difficult to achieve in the early stages of training, we propose an adaptive control factor ($\lambda_{\text{ctrl}}$) strategy updated according to training performance (see Appendix A).
Improved Clustering Our loss design not only achieves dimensionality reduction and clustering but also enhances clustering properties, making the states within each cluster more compact (smaller intra‑cluster distances) and the cluster boundaries more separable. This is crucial for clearly distinguishing the semantics of states at the cluster boundaries, further enhancing the model’s interpretability. Because of the stop-gradient in $\mathcal{L}^{\prime}_{\text{VQ-VAE}}$, it does not directly pull FDR features toward their nearest codebook embeddings. However, when the FDR features become denser during joint training with $\mathcal{L}_{\text{FDR}}$, $\mathcal{L}^{\prime}_{\text{VQ-VAE}}$—and thus $\mathcal{L}_{\text{total}}$—decreases, yielding tighter clusters. Moreover, since $\mathcal{L}_{\text{FDR}}$ aligns the affinity matrices $p$ and $q$, this densification in the low-dimensional FDR space is reflected and propagates into the high-dimensional state features. We demonstrate the improved clustering in § 4 and provide more evidence of this enhanced clustering property and the intrinsic nature of semantic clustering in DRL in § C.2.
@@
2
31exfor each training iteration $i=1,2,\dots$ do
@@
7
       $\mathcal{L}_{\text{PPO}}\leftarrow\mathcal{L}_{\text{policy}}+w_{\text{value}}\mathcal{L}_{\text{value}}+w_{\text{entropy}}\mathcal{L}_{\text{entropy}}$;
@@
       $\mathcal{L}_{\text{SCM}}\leftarrow w_{\text{FDR}}\mathcal{L}_{\text{FDR}}+w_{\text{VQ-VAE}}\mathcal{L}^{\prime}_{\text{VQ-VAE}}$;
@@
       $\mathcal{L}_{\text{total}}\leftarrow\mathcal{L}_{\text{PPO}}+\lambda_{\text{ctrl}}\mathcal{L}_{\text{SCM}}$;
@@
Advantages of Online Training Online training offers several advantages: (i) it enhances clustering by incorporating the training of $\mathcal{L}_{\text{total}}$, (ii) training the VQ code $k$ with a latent-conditioned policy $\pi\left(a|\mathbf{s},k\right)$ (where $a$ is the action) supports extension to downstream tasks, such as macro action selection in hierarchical learning, and (iii) it improves memory efficiency by eliminating the need to store a large number of states during model training.
@@ 4 Simulations
In this work, we primarily study the intrinsic characteristics and black-box decision-making of DRL, and address the instability of t-SNE visualizations used in prior studies. Therefore, this section aims to: (i) compare t-SNE to validate the stability and effectiveness of the proposed clustering method, (ii) assess the semantic clustering properties of DRL to improve interpretability, and (iii) introduce new methods to analyze policies and internal model characteristics, identifying issues in DRL decision-making. The integration of our module has minimal impact on performance (see § C.1 and Appendix D).
@@ 4.1 Clustering Effectiveness Evaluation
We demonstrate the clustering effectiveness of our proposed approach using the CoinRun game from Procgen as an example. Similar results can be easily extended to other games using the code and checkpoints provided in the supplementary material. We use a trained model to collect states, where the agent selects actions randomly with a probability of 0.2 and follows the trained policy with a probability of 0.8 to ensure diverse state coverage. States are sampled with a probability of 0.8, and 64 parallel environments collect states over 500 steps, resulting in $\approx$25,000 states for visualization. Note that the cluster colors (indices) in the t-SNE plots are assigned by our method and are used solely to facilitate comparison of spatial relationships.
Cluster Separation and Improved Clustering The t-SNE visualization of PPO (2(a)), spreads features across the space without forming clear clusters, limiting its utility for clustering analysis and requiring detailed manual examination of certain areas, as in previous studies. In contrast, the t-SNE visualization of our method (2(b)), reveals numerous distinct, small clusters. States within each of these clusters originate from the same semantic group identified by our method. This dispersion into multiple smaller clusters is due to t-SNE’s focus on local structures and its tendency to avoid crowding, causing complete semantic clusters to scatter. The visualization in the FDR space (2(d)), displays clear and separate complete clusters, which are identified by VQ codes. § C.2 presents a stop-gradient ablation that disables our proposed module while keeping the backbone and training protocol fixed; visualizations show fuzzier, less separable clusters than the full model, further supporting our gains in sharpness and coherence.
@@
Sensitivity to Random Seed While the t-SNE representation is sensitive to randomness, as demonstrated by the significant difference between Figures 2(c) and 2(f), the FDR space’s mapping remains unchanged even when the random seed is altered, maintaining the distribution in 2(e). t-SNE’s randomness primarily stems from its random initialization and non-convex optimization process, leading to significantly different visualizations with different random seeds. In contrast, our model produces a stable feature mapping after training, which does not vary with random seeds.